A student taking a quiz finds on a reference sheet the two equations f = 1 / T and =square root of (1/T).
She has forgotten what T represents in each question.
Use dimensional analysis to determine the units required for T in each question.
Explain how you can identify the physical quantity each T represents from the units.
1
Expert's answer
2020-10-07T07:22:48-0400

The reciprocal of period of oscillation is the frequency

"f =\dfrac{1}{T}"

f is frequency and T is period

Rearranging the equation for T

"T =\dfrac{1}{f}"

using the dimensional analysis to write units for period

"[T] =\dfrac{1}{[f]}"

substituting dimension formula T-1 for f

"[T] =\dfrac{1}{[\left (T^{\smash{-1}}\right)]}"

=T

where T is time

"\therefore" units for T are seconds

The expression for velocity of wave is,

"v=\sqrt{\dfrac{T}{\mu}}"

where T is tension in the string , and "\mu" is mass per unit length.

Rearranging the equation for T

"T=\mu*v^2"

using dimensional analysis to write units for the period

"[T]=[\mu]*[v^2]"

substituting dimensional formula "\frac{M}{L}" for "\mu" and "\frac{L^2}{T^2}" for v.

"[T]=\frac{M}{L} \frac{L^2}{T^2}"

"=\frac{ML}{T^2}"

Here M represents mass, L represents length, and T is time

Therefore, the units for tension is kg.m/s2 or newtons


b). In the first part, T represents period of time . in the second part of the solution . T represents the force of tension.